# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ A.7 Proof of Theorem 4
    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\sigma_{\gamma}^{2}C_{f}^{2}\mathrm{len}_{\gamma}^{2}C_{\gamma,f}^{2}}{\mathrm{reach}_{\gamma}^{2}C_{f}^{\prime 2}}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}$
    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\mathrm{len}_{\gamma}R_{0}^{2}}{\sigma_{\gamma}^{2}}\right)^{2(s\wedge 1)}(d^{2}\log d)^{s\wedge 1}C_{\gamma,f}^{2(s\wedge 1)}\frac{(\log n)^{7(s\wedge 1)}}{n^{4(s\wedge 1)}}$
    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2s}C_{\gamma,f}^{2s}\frac{(\log n)^{3s}}{n^{2s}}$
@@ A.8 Proof of Theorem 3
    $n^{\frac{1}{2s+1}}M^{*}\text{ where }M^{*}=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}$
@@
    $\frac{C_{f}^{\frac{2s}{2s+1}}}{C_{f}^{\prime\frac{2s}{2s+1}}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}n^{-\frac{2s}{2s+1}}{\log n}$ (8)
@@
    $\epsilon_{0}^{2}=\max\left(\frac{\left[f\right]_{\mathcal{C}^{s}}^{2}\sigma_{\gamma}^{2s}C_{f}^{2s}\max(\sigma_{\zeta},\omega_{f})^{2s}}{\mathrm{reach}_{\gamma}^{2s}},\frac{C_{f}^{\frac{2s}{2s+1}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}\log n}{C_{f}^{\prime\frac{2s}{2s+1}}n^{\frac{2s}{2s+1}}}\right)$
@@ A.10 Proof of Proposition 14 and Theorem 5
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (SC), and ($\bm{\omega_{f}}$) hold true. Let ${\mu_{l,h}^{b}}$ be the mean of $h$-th slice and ${v_{l,h}^{b}}$ be the significant vector of $h$-th slice. Then, for every $l$ such that $|{R_{l,h}^{b}}|\asymp\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon\in(0,1)$ and $\tau\geqslant 1$, on each slice
@@
      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+n^{-\tau}$
@@
      $\mathbb{P}\left\{\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|>\epsilon\ \big{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
@@
      $\qquad\qquad\lesssim d\exp\left(-c\frac{u^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{2}+uC_{f}^{2}|{R_{l,h}^{b}}|^{2}R_{0}^{2}d}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
@@
    $\mathbb{P}\left(\max\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|,\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\right)\gtrsim\beta C_{f}^{2}|{R_{l,h}^{b}}|^{2}\right)\lesssim d\exp\left(-\frac{c\beta^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)$
@@
    $\bm{\mathfrak{S}}(\epsilon,\beta,u)=\left\{\begin{split}&\text{ For all }h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\frac{R_{0}\sqrt{d}}{2},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon\\ &\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2},\max\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2},\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\right)\lesssim\beta^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\end{split}\right\}$
@@
    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|$
@@
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim C_{f}|{R_{l,h}^{b}}|\ \ \text{ for any }|k|\geqslant 2$
@@
    $\lesssim$ $C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}+C_{f}|{R_{l,h}^{b}}|\beta\sqrt{{\mathrm{dist}}(x,h)}+u^{2}R_{0}^{2}d+\epsilon^{2}R_{0}^{2}d+C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}\epsilon^{2}+\frac{C_{f}^{4}|{R_{l,h}^{b}}|^{4}}{R_{0}^{2}d}\epsilon^{2}\lesssim C_{f}^{2}|{R_{l,h}^{b}}|^{2}$
@@
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}$
@@ Appendix B Technical Results
    $\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>t\right\}\leqslant 2\exp\left\{-\frac{nt^{2}/2}{\rho(E)+t/3}\right\}$
@@
    $\mathbb{P}\left\{\hat{\rho}(E)\not\in\left[\frac{1}{2}\rho(E),\frac{3}{2}\rho(E)\right]\right\}\leqslant\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>\frac{1}{2}\rho(E)\right\}\leqslant 2\exp(-\frac{3}{28}n\rho(E))$
@@
    Assume (X subG), ($\bm{\zeta}$ subG), and (P). Suppose function $\widetilde{f}\in\mathcal{C}^{s}$ with $s\in[\frac{1}{2},1]$ is defined on interval $I$ with diameter $r\geqslant 1$. Let integer $m=\lfloor s\rfloor$ be the largest integer smaller than or equal to $s$, so $m=0$ if $s\in[\frac{1}{2},1)$ and $m=1$ if $s=1$. Given $j\geqslant 1$, construct $\{I_{j,k}\}_{k=1}^{j}$ the uniform partition of $I$ into $j$ intervals. Let ${\widehat{v}}\in\mathbb{S}^{d-1}$ be an estimator of $v\in\mathbb{S}^{d-1}$ such that $\langle{\widehat{v}},v\rangle\geqslant 0$. Given data $\{(X_{i}^{\prime},Y_{i}^{\prime})\}_{i=1}^{n^{\prime}}$. For $u\in\{v,{\widehat{v}}\}$, denote $n_{j,k|u}=\#\{X_{i}^{\prime}:\langle u,X_{i}^{\prime}\rangle\in I_{j,k}\}$. For each $k\in\mathcal{K}_{j|u}:=\{k:n_{j,k|u}\geqslant n^{\prime}/j\}$, compute
    ${\widehat{f}_{j,k|u}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{i\leqslant n^{\prime}}\left|Y_{i}-p(\langle u,X_{i}^{\prime}\rangle)\right|^{2}\mathbbm{1}_{I_{j,k}}(\langle u,X_{i}^{\prime}\rangle)$
    and compute the piecewise constant $(s<1)$ or linear $(s=1)$ estimator of $\widetilde{f}$ at scale $j$ conditioned on $u$ and truncated at $\left|\widetilde{f}\right|_{L^{\infty}}$
    ${\widehat{f}_{j|u}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|u}}(r)\mathbbm{1}_{I_{j,k}}(r)$
@@
    $\mathbb{E}_{X}\left[\left|{\widehat{f}_{j,k|v}}(\langle v,X\rangle)-{\widehat{f}_{j,k|{\widehat{v}}}}(\langle{\widehat{v}},X\rangle)\right|^{2}\mathbbm{1}_{B(0,r)}(X)\right]\lesssim\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{\frac{2}{2-s}}\left\|{\widehat{v}}-v\right\|^{\frac{2}{2-s}}+\epsilon_{0}^{2}$
    with probability higher than $1-C\#\mathcal{K}_{j}\exp\left(-\frac{cn^{\prime}\epsilon_{0}^{2}}{\#\mathcal{K}_{j}\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{2s}}\right)$

### v2

@@ A.7 Proof of Theorem 4
    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\sigma_{\gamma}^{2}C_{f}^{2}\mathrm{len}_{\gamma}^{2}C_{\gamma,f}^{2}}{\mathrm{reach}_{\gamma}^{2}C_{f}^{\prime 2}}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}\ \,;$
    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\mathrm{len}_{\gamma}R_{0}^{2}}{\sigma_{\gamma}^{2}}\right)^{2(s\wedge 1)}(d^{2}\log d)^{s\wedge 1}C_{\gamma,f}^{2(s\wedge 1)}\frac{(\log n)^{7(s\wedge 1)}}{n^{4(s\wedge 1)}}\ \,;$
    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2s}C_{\gamma,f}^{2s}\frac{(\log n)^{3s}}{n^{2s}}\ \,;$
@@ A.8 Proof of Theorem 3
    $n^{\frac{1}{2s+1}}M^{*}\text{ where }M^{*}=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}\ \,.$
@@
    $\frac{C_{f}^{\frac{2s}{2s+1}}}{C_{f}^{\prime\frac{2s}{2s+1}}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}n^{-\frac{2s}{2s+1}}{\log n}\ \,.$ (8)
@@
    $\epsilon_{0}^{2}=\max\left(\frac{\left[f\right]_{\mathcal{C}^{s}}^{2}\sigma_{\gamma}^{2s}C_{f}^{2s}\max(\sigma_{\zeta},\omega_{f})^{2s}}{\mathrm{reach}_{\gamma}^{2s}},\frac{C_{f}^{\frac{2s}{2s+1}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}\log n}{C_{f}^{\prime\frac{2s}{2s+1}}n^{\frac{2s}{2s+1}}}\right)\ \,.$
@@ A.10 Proof of Proposition 14 and Theorem 5
[…]
@@
      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+n^{-\tau}\ \,.$
@@
      $\mathbb{P}\left\{\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|>\epsilon\ \big{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\ \,.$
@@
      $\qquad\qquad\lesssim d\exp\left(-c\frac{u^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{2}+uC_{f}^{2}|{R_{l,h}^{b}}|^{2}R_{0}^{2}d}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\ \,.$
@@
    $\mathbb{P}\left(\max\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|,\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\right)\gtrsim\beta C_{f}^{2}|{R_{l,h}^{b}}|^{2}\right)\lesssim d\exp\left(-\frac{c\beta^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)\ \,.$
@@
    $\bm{\mathfrak{S}}(\epsilon,\beta,u)=\left\{\begin{split}&\text{ For all }h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\frac{R_{0}\sqrt{d}}{2},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon\\ &\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2},\max\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2},\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\right)\lesssim\beta^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\end{split}\right\}\ \,.$
@@
    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|\ \,.$
@@
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim C_{f}|{R_{l,h}^{b}}|\ \ \text{ for any }|k|\geqslant 2\ \,.$
@@
    $\lesssim$ $C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}+C_{f}|{R_{l,h}^{b}}|\beta\sqrt{{\mathrm{dist}}(x,h)}+u^{2}R_{0}^{2}d+\epsilon^{2}R_{0}^{2}d+C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}\epsilon^{2}+\frac{C_{f}^{4}|{R_{l,h}^{b}}|^{4}}{R_{0}^{2}d}\epsilon^{2}\lesssim C_{f}^{2}|{R_{l,h}^{b}}|^{2}\ \,,$
@@
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}\ \,.$
@@ Appendix B Technical Results
    $\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>t\right\}\leqslant 2\exp\left\{-\frac{nt^{2}/2}{\rho(E)+t/3}\right\}\ \,.$
@@
    $\mathbb{P}\left\{\hat{\rho}(E)\not\in\left[\frac{1}{2}\rho(E),\frac{3}{2}\rho(E)\right]\right\}\leqslant\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>\frac{1}{2}\rho(E)\right\}\leqslant 2\exp(-\frac{3}{28}n\rho(E))\ \,.$
@@
[…]
    ${\widehat{f}_{j,k|u}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{i\leqslant n^{\prime}}\left|Y_{i}-p(\langle u,X_{i}^{\prime}\rangle)\right|^{2}\mathbbm{1}_{I_{j,k}}(\langle u,X_{i}^{\prime}\rangle)\ \,,$
    and compute the piecewise constant $(s<1)$ or linear $(s=1)$ estimator of $\widetilde{f}$ at scale $j$ conditioned on $u$ and truncated at $\left|\widetilde{f}\right|_{L^{\infty}}$:
    ${\widehat{f}_{j|u}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|u}}(r)\mathbbm{1}_{I_{j,k}}(r)\ \,.$
@@
    $\mathbb{E}_{X}\left[\left|{\widehat{f}_{j,k|v}}(\langle{\widehat{v}},X\rangle)-{\widehat{f}_{j,k|{\widehat{v}}}}(\langle{\widehat{v}},X\rangle)\right|^{2}\mathbbm{1}_{B(0,r)}(X)\right]\lesssim\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{\frac{2}{2-s}}\left\|{\widehat{v}}-v\right\|^{\frac{2}{2-s}}+\epsilon_{0}^{2}\ \,.$
    with probability higher than $1-C\#\mathcal{K}_{j}\exp\left(-\frac{cn^{\prime}\epsilon_{0}^{2}}{\#\mathcal{K}_{j}\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{2s}}\right)$.
    Note that the assumption ($\bm{\text{P}_{\gamma}}$) uses the push-forward of the distribution of $X$ along the map $x\mapsto\Pi_{\gamma}(x)$, while in this Lemma we use the push-forward along the orthogonal projection onto the tangent vector at $\Pi_{\gamma}(x)$. This second measure however does inherit a lower-bound on its density by the regularity of the curve and of the two maps, the second one being a local linearization of the first one.